I would gladly appreciate if i could get some help.
Problem Set 1A/1: Trade Name Do this problem set if your surmame (family name) begins R-z. The small nations of Arboc and Arbez quantity of each good is shown by the production possibili nations trade in à world containing many small nations. Costs are constant can produce two goods, wheat and coffee. The maximum ty information below. Assume the two Arboc Arbez wheat wheat 50 and 0 and 100 30 and 0 0 and 0 1. What is the opportunity cost of one unit of wheat in Arbed 100/60-2 2 What is the opportunity cost of one unit of wheat in Arbez?90/Io 3 3. Which country has a comparative advantage in the production of wheat? 4. What is the opportunity cost of one unit of coffee in Arboc? 5. What is the opportunity cost of one unit of coffee in Arbez? 6. Which country has a comparative advantage in the production of coffee? 7. The terms of trade are 1 unit of coffee per unit of wheat. Arboc should produce which good? & The terms of trade are 2.5 units of coffee per unit of wheat. Arboc should produce which good? Arbez should produce which good? Arbez should produce which good? The terms of trade are 3 units of coffee per unit of wheat. Which country 9. The terms of trade are 4 units of coffee per unit of wheat. 10. could not gain from trade?
Answer #1

According to the question,

Wheat

coffee

Arboc

50

100

Arbez

30

90

Opportunity cost calculated:

Wheat

coffee

Arboc

100/50=2

50/100 = 1/2

Arbez

90/30 =3

30/90=1/3

1. The opportunity cost of producing 1 unit of wheat in Arboc = 2 units of coffee

2. The opportunity cost of producing 1 unit of wheat in Arbez= 3 units of coffee

3. Comparing the opportunity cost (2<3) Arboc has the comparative advantage of producing wheat as it has a lower opportunity cost in producing it.

4. The opportunity cost of producing 1 unit of coffee in Arboc = 1/2 units of coffee

5. The opportunity cost of producing 1 unit of coffee in Arbez= 1/3 units of coffee

6. Comparing the opportunity cost (1/2>1/3) Arbez has the comparative advantage of producing coffee.

7. The terms of trade are 1 unit of coffee per unit of wheat. Arboc should produce coffee.

8. The terms of trade are 1 unit of coffee per unit of wheat. Arboc should produce coffee. Arbez should also produce coffee.

9. The terms of trade are 4 units of coffee per unit of wheat. Arbez should also produce coffee.

10. The terms of trade are 3 units of coffee per unit of wheat. Arbez should also produce coffee.